# Game-based Platforms for Artificial Intelligence Research

### v3

@@ III-B Video Games
This section categorises the video games into grid-based 2D, pixel-based 2D and 3D according to the simulated physical space with which avatar(s) interact. Grid-based games are usually constructed by some preset tiles where players act with a set of discrete actions. Unlike grid-based games, pixel-based games feature continuous state space with either continuous or discrete action space. The difference in the state representation results in different feature extraction between grid-based and pixel-based games. The state in grid-based games can be easily converted to a low-dimensional matrix since the number of tiles is usually countable. On the other hand, feature extraction for pixel-based games is usually applied to states represented by high-dimensional images. 3D games extend 2D games with an additional height dimension, which not only enriches the visual and interactive depth of games but also introduces significant challenges in representing the state effectively. Besides, the action spaces of 3D games are usually more complicated or larger than those of 2D games.
@@
III-B1 Grid-based 2D Games
@@
Classic single-player and multi-player competitive games have been implemented using grid-based maps. SimCity [41] is an open-ended city-building 2D video game with an open-source version provided for reinforcement learning research. Nethack [42] is a single-agent roguelike game with procedurally generated levels. Agents meet the different stages at each episode, which drives AI’s abilities for long-term exploration and generalisation. Compared with existing single-player games, Nethack provides a faster simulator with massive tasks and game dynamics. Multi-agent games involve multiple independent decision agents, oriented towards cooperative and competitive behaviours. Snake AI [43] allows two snakes to fight for an apple. Since the apple is placed randomly in maps, the game can sometimes be unbalanced.
@@
Modern strategy games incorporate digital elements to achieve complex and diverse mechanisms. microRTS [44] is a lightweight, grid-based real-time strategy (RTS) game framework that has fewer rendering sprites and focuses on the game’s mechanics, in which units are represented by different colours and shapes. Compared with StarCraft II, microRTS is easier to manipulate, so that researchers can focus more on the design of AI ideas and algorithms. As an implementation of The Battle of Polytopi, Tribe [45] involves multiple multi-agent scenarios as a turn-based strategy game framework, in which a forward model is supported for decision. Partial observability mode is supported by the framework, where the unexplored area is covered by the fog of war.
@@
The work of Carroll et al. [46] provides a simple implementation of the commercial game Overcooked! for human-AI coordination, in which agents and players work together to cook meals. Agents and players need to move discretely and avoid colliding with each other, which introduces a new challenge of making motion coordination strategies. However, this implementation [46] encodes the game map as a 2D array, which can hardly express the complete information of Overcooked!. For instance, the representation of the game level can not distinguish whether an onion is on a table or the floor. Besides, the ingredients are fixed in the implementation. It’s worth mentioning that the work of [47] also introduces a game engine similar to Overcooked!, called CuisineWorld, to evaluate multi-agent cooperation performance. The platform has been tested by LLM agents [47]. Though CuisineWorld is a 3D game engine, to our best knowledge, 2D game observations are used in [47]. Instead of traditional representation based on visual information, CuisineWorld inverts the game state with text-based representation and the action space consists of several instructions like moving and putting. Compared to the implementation by [46], CuisineWorld is more suitable for research on LLMs. Pommerman [48] is a multi-player competitive game for at least four players, who start with a limited number of bombs. Players move in a grid-based map and plant the bombs to destroy their opponents. Pommerman allows multiple players to cooperate and compete as a team on the battlefield. Neural Massive Multiplayer Online (Neural MMO) [49] considers a large multi-agent scenario where at most 1024 agents learn to cooperate and compete to survive in the grid-based world at the same time.
@@
III-B2 Pixel-based 2D Games
Pixel-based games support more complex game mechanisms with continuous state space, while new challenges are introduced.
Mario AI framework [50] is based on the game Super Mario Bros., where players control Mario to run and jump through some barriers in a partially observable environment to arrive at the endpoint. However, the provided levels are simple and a limited number of levels are available, planning algorithms such as $*{}^{*}$ can quickly solve it without leveraging sophisticated AI approaches. MsPacMan-vs-Ghosts [51] is the implementation of Ms. Pac-Man arcade game, in which the player controls Ms. Pac-Man to collect pills while avoiding being hunted by four ghosts. Both full and partial observability are supported in MsPacMan-vs-Ghosts. AIBIRDS [52] is the platform with the popular puzzle game Angry Birds. Agents are required to shoot pigs to destroy the building made of blocks of different types and shapes. XFC [53] simulates the game Asteroid Smashers, in which AI controls a spacecraft to avoid asteroids and destroy pesky rocks. XFC designs a specific agent architecture, which is suitable for research on explainable AI and fuzzy AI.
Geometry Friends [54] provides a cooperative scenario that two agents aim to solve puzzles, restricted by real physics conditions such as friction and gravity. A simple implementation of Planet Wars, named AIGD2 [55], brings a new challenge with extra-long skill-depth to RTS games. FightingICE [56] considers a 2D video fighting game that AI agents fight against opponents according to sound. “Blind" agents face challenges such as partial observability and insufficient training.
Honor of Kings Arena, Dota 2 and StarCraft II [6] are three popular commercial games with publicly available APIs that can be used for AI research. Honor of Kings Arena [57] provides an interface to interact with Honor of Kings 1v1, in which agents learn to fight against diverse opponents. An API of Dota 2 written with Lua language is available for investigating multi-agent RL [7]. Fig. 2 presents the heroes controlled by OpenAI Five [7]. An interface connecting to StarCraft II [6] is also released for research in multi-agent RL and map design. Source codes of commercial games are often inaccessible, which means AI researchers can hardly edit commercial games to meet some particular requirements for research, such as creating an abstract, faster forward model of the game, even though their APIs are sometimes available.
@@ III-B2 Pixel-based 2D Games
ViZDoom [58] is a highly customisable doom-based environment with partial observability. Raw visual screens from the first-person perspective are received by the agents. ViZDoom is lightweight and fast. Its integrated editor allows users to manually create diverse task suits. ViZDoom is suitable as the test-bed for solving visual RL problems.
@@
MineRL [59] is based on Minecraft, where a large scale human dataset is provided. The work of Baker et al. [61] presents a specific platform for Hide and Seek, which can be used for training and understanding multi-agent cooperative and competitive AI. MineDojo [60] provides a Gym-like environment based on Minecraft. Agents are supported to directly access the game for multiple tasks including material mining, crafting and building. Besides, a massive multimodal dataset, captured from video websites, gameplays, wiki pages and Reddit posts is provided.
@@
Figure 4: Summary of games and platforms used for studing AI for game design, created by authors using screenshots of the games and platforms, as well as figures from their corresponding papers, reviewed in Section IV.
III-C Platforms
@@
Classic and modern board, card and puzzle games are often integrated into a single platform as the AI test-bed. For example, General Game Playing (GGP[62] is designed to evaluate AI’s ability in playing different kinds of tabletop games, where rules and states of games are defined with game description language, and has been widely used in AI research and education. Hundreds of single- and multi-player games, such as Go, Tic-Tac-Toe, reversi and minesweeper, are available in GGP. Botzone [63] focuses more on large multi-player games and is famous for its Mahjong AI Competition. Botzone also support 2- and 6-person Texas hold’em. Besides some popular classic games, OpenSpiel [64] provides Poker and Bridge, as well as some classic algorithms and visual tools. Ludii [88] is another general game-playing system based on similar principles to Ludi [89] but with different mechanisms. It includes various types of games such as board games, boardless games, stacking games, single- and multi-player games. Ludii uses ludemes as game description grammar. The description language is easy to read and understand according to [88]. Different from GGP and Botzone, the Tabletop Games Framework (TAG) [66] is designed to study general AI in modern tabletop games and has implemented more than 20 games, including Uno, Love Letter and Virus!. An extra analysis tool is equipped in TAG for studying action space, branching factors, hidden information, and so on. Pgx [67] is a environment of board games for RL research. It is written in JAX and can do parallel simulation via GPU/TPU. Pgx can hold single-/multi-agent board games and Atari-like environments including Shogi, Go, Chess, and Backgammon. According to [67], Pgx can perform 10-100 times higher throughput compared to other Python libraries.
@@
OpenAI Gym [68] and PettingZoo [69] offer a number of grid-based and pixel-based 2D games. OpenAI Gym [68] provides an API standard for RL research and contains massive games, such as Atari, as well as a wide range of third-party environments. However, Gym typically focuses on the single-agent setting. To conduct research in multi-agent RL, PettingZoo [69] extends the API standard of Gym and integrates some famous environments like multi-player Atari games, classic card/board games and multi-agent particle environment [90]. As providing the standard API, Gym and PettingZoo are now the most used and compatible platforms, which enhance the evaluation and reproduction in the RL community with a higher speed.
@@
Minigrid [70] is a single-agent platform compatible with OpenAI Gym. The included games are simple in rendering content but can implement complex mechanisms. General Video Games AI (GVGAI) [71] platform and GVGAI gym [72] offer a larger number of grid-based single- and two-player games, and has been used in its single-player planing, two-player planing, and single-player learning competitions. Users can create new games and levels in GVGAI in an intuitive way using Video Game Description Language (VGDL) [91]. Stratega [73] evaluates the generalisation of strategy agents with multiple games and levels with an $n$-player turn-based setting. Profiling and logging tools are integrated for analysis. Usually, grid-based platforms are highly scalable in terms of game mechanisms and levels, so researchers can easily design specific tasks to address their research problems with the help of grid-based game platforms. Griddly [74] particularly provides grid-based real-time strategy games with various numbers of players, with a faster running speed and lower memory than Minigrid and GVGAI. Melting Pot [75] is a platform with more than 50 diverse games for evaluating the generalisation of multi-agent RL.
@@
DeepMind Lab [79] acts as a 3D learning environment which is based on some open source software such as ioquake3. It provides a set of challenging 3D navigation and puzzle-solving tasks for agents to learn. The primary purpose of DeepMind Lab is to act as a working platform for results in AI, especially deep reinforcement learning. ML-agents [80] is an open-source platform that allows to create new games in a simpler and more flexible manner with the help of the Unity3D toolkit. MiniWorld [81] provides a more accessible way to construct 3D environments for reinforcement learning and robotics research. It is qualified to generate and simulate environments with rooms, doors, mazes and other objects. MiniWorld is simpler than DeepMind Lab and ViZDoom in the view of engineering, so beginners can get started quickly without mastering Unity3D.
III-D Guidance for Playing Games
According to the aim, number of agents, observability, language and competition, games and platforms reviewed in this work are categorised in Tabs. I and II. Those games and platforms are often developed for specific research purposes.
Card/board-related games and platforms, including Blood bowl [29], Sabberstone [34] and Botzone [63], are suitable for research on tree search, causality and competitive scenarios in multi-agent systems. GVGAI [71, 72, 92], Stratega [73] and Procgen [77] are developed for the purpose of evaluating the generalisation of AI agents. Both GVGAI and Stratega can be used for research on planning and learning, while Procgen specially focus on RL.
@@ III-D Guidance for Playing Games
Besides, the integrated tools differ between games and platforms. For example. in GVGAI, users can easily construct their games and levels via VGDL. Pommerman [48] supports highly self-defined maps. However, game levels in Procgen can only be generated randomly. It is recommended to get started from Gym [68] and PettingZoo [69] for single- and multi-agent RL research since they are the most common platforms. Some platforms such as DeepMind Lab [79] and ML-agent [80] usually are equipped with some visual tools and loggers to analyse the behaviour of agents, but the requirement to learn about the game engines like Unity is crucial.
@@
IV Designing Games by AI
TABLE III: 25 environments for designing games by AI. Although Mario-AI, HOG2, PhyDSL, StepMania, OSU! and MiniDungeons are available before 2016, they have been used for testing and evaluating the procedurally generated SMB levels in recent years.
@@ IV Designing Games by AI
Game Tuning (Section IV-A) Parameter FightingICE Java [93]
@@
Content Generation (Section IV-B) 2D level Mario AI Java [50]
@@
Science Birds Java [97]
Morai-Maker-Engine C# [98]
@@
Other (Section IV-C) Code PhyDSL Java [110]
@@
In recent years, game-based platforms have also been used to create new tasks for AI and to study AI creativity, in particular, the ability of creating new contents. Tab. III and Fig. 4 summarise the games and platforms used for game design. Those platforms provide diverse and representative test-beds for discovering the potentials in AI-involved game design.
IV-A Game Tuning

### v4

@@ 3.2 Video Games
This section categorises video games into grid-based 2D, pixel-based 2D and 3D according to the simulated physical space with which avatar(s) interact. Grid-based games are usually constructed by some preset tiles where players act with a set of discrete actions. Unlike grid-based games, pixel-based games feature continuous state space with either continuous or discrete action space. The difference in the state representation results in different feature extraction between grid-based and pixel-based games. The state in grid-based games can be easily converted to a low-dimensional matrix since the number of tiles is usually countable. On the other hand, feature extraction for pixel-based games is usually applied to states represented by high-dimensional images. 3D games extend 2D games with an additional height dimension, which not only enriches the visual and interactive depth of games but also introduces significant challenges in representing the state effectively. Besides, the action spaces of 3D games are usually more complicated or larger than those of 2D games.
@@
3.2.1 Grid-based 2D Games
@@
Classic single-player and multi-player competitive games have been implemented using grid-based maps. Snake AI [63] allows two snakes to fight for an apple. Since the apple is placed randomly in maps, the game can sometimes be unbalanced. SimCity [64] is an open-ended city-building 2D video game with an open-source version provided for RL research. Nethack [65] is a single-agent roguelike game with procedurally generated levels. Agents meet the different scenarios at each episode, which drives AI’s abilities for long-term exploration and generalisation. Compared with existing single-player games, Nethack provides a faster simulator with massive tasks and game dynamics. Multi-agent games involve multiple independent decision agents, oriented towards cooperative and competitive behaviours. Pommerman [66] is a multi-player competitive game for at least four players, who start with a limited number of bombs. Players move in a grid-based map and plant the bombs to destroy their opponents. Pommerman allows multiple players to cooperate and compete as a team on the battlefield. Neural Massive Multiplayer Online (Neural MMO) [67] considers a large multi-agent scenario where at most 1024 agents learn to cooperate and compete to survive in the grid-based world at the same time.
@@
Modern strategy games incorporate digital elements to achieve complex and diverse mechanisms. microRTS [68] is a lightweight, grid-based RTS game framework that has fewer rendering sprites and focuses on the game’s mechanics, in which units are represented by different colours and shapes. Compared with StarCraft II, microRTS is easier to manipulate, so that researchers can focus more on the design of AI ideas and algorithms. As an implementation of The Battle of Polytopi, Tribe [69] involves multiple multi-agent scenarios as a turn-based strategy game framework, in which a forward model is supported for decision. Partial observability mode is supported by the framework, where the unexplored area is covered by the fog of war.
@@
The work of Carroll et al. [70] provides a simple implementation of the commercial game Overcooked! for human-AI coordination, in which agents and players work together to cook meals. Agents and players need to move discretely and avoid colliding with each other, which introduces a new challenge of making motion coordination strategies. However, this implementation [70] encodes the game map as a 2D array, which can hardly express the complete information of Overcooked!. For instance, the representation of the game level can not distinguish whether an onion is on a table or the floor. Besides, the ingredients are fixed in the implementation. It’s worth mentioning that the work of [71] also introduces a game engine similar to Overcooked!, called CuisineWorld, to evaluate multi-agent cooperation performance. The platform has been tested by LLM agents [71]. Though CuisineWorld is a 3D game engine, to our best knowledge, 2D game observations are used in [71]. Instead of traditional representation based on visual information, CuisineWorld inverts the game state with text-based representation and the action space consists of several instructions like moving and putting. Compared to the implementation by [70], CuisineWorld is more suitable for research on LLMs.
@@
3.2.2 Pixel-based 2D Games
Pixel-based games support more complex game mechanisms with continuous state space, while new challenges are introduced. Mario AI framework [72] is based on the game Super Mario Bros., where players control Mario to run and jump through some barriers in a partially observable environment to arrive at the endpoint. However, the provided levels are simple and a limited number of levels are available, planning algorithms such as A can quickly solve it without leveraging sophisticated AI approaches. MsPacMan-vs-Ghosts [73] is the implementation of Ms. Pac-Man arcade game, in which the player controls Ms. Pac-Man to collect pills while avoiding being hunted by four ghosts. Both full and partial observability are supported in MsPacMan-vs-Ghosts. Geometry Friends [74] provides a cooperative scenario that two agents aim to solve puzzles together, restricted by real physics conditions such as friction and gravity. A simple implementation of Planet Wars, named AIGD2 [75], brings a new challenge with extra-long skill-depth to RTS games. XFC [76] simulates the game Asteroid Smashers, in which AI controls a spacecraft to avoid asteroids and destroy pesky rocks. XFC designs a specific agent architecture, which is suitable for research on explainable AI and fuzzy AI. FightingICE [77] considers a 2D fighting video game that AI agents fight against opponents according to sound. “Blind" agents face challenges such as partial observability and insufficient training. AIBIRDS [78] is the platform with the popular puzzle game Angry Birds. Agents are required to shoot pigs to destroy the building made of blocks of different types and shapes.
nsh [79] is a fight game simulator for the commercial massive multiplayer online (MMO) game Justice Online. Renders and memory usage are largely reduced in nsh, while game rules are close to the ones in the original game. Honor of Kings Arena [80] provides an interface to interact with Honor of Kings 1v1, in which agents learn to fight against diverse opponents. An API of Dota 2 written with Lua language is available for investigating multi-agent RL [16]. Fig. 2 presents the heroes controlled by OpenAI Five [16]. Dunk City Dynasty simulates a 3v3 basketball commercial game with the same name [81]. A dataset with 400GB human data and records is also provided. An interface connecting to StarCraft II [15] is also released for research in multi-agent RL and map design. Source codes of commercial games are often inaccessible, which means AI researchers can hardly edit commercial games to meet some particular requirements for research, such as creating an abstract, faster forward model of the game, even though their APIs are sometimes available.
@@ 3.2.2 Pixel-based 2D Games
MineRL [82] is based on Minecraft, where a large-scale human dataset is provided. MineDojo [83] provides a Gym-like environment based on Minecraft. Agents are supported to directly access the game for multiple tasks including material mining, crafting and building. Besides, a massive multimodal dataset, captured from video websites, gameplays, wiki pages and Reddit posts is provided. Google Research Football [84] provides a real-world football challenge with a physical 3D engine, in which multiple tasks with variable difficulty are available. ViZDoom [85] is a highly customisable doom-based environment with partial observability. Raw visual screens from the first-person perspective are received by the agents. ViZDoom is lightweight and fast. Its integrated editor allows users to manually create diverse task suits. ViZDoom is suitable as the test-bed for visual RL research. The work of Baker et al. [86] presents a specific platform for Hide and Seek, which can be used for training and understanding multi-agent cooperative and competitive AI.
@@
Figure 4: Summary of games and platforms used for studing AI for game design, created by authors using screenshots of the games and platforms, as well as figures from their corresponding papers, reviewed in Section 4.
3.3 Platforms
@@
Classic and modern board, card and puzzle games are often integrated into a single platform as the AI test-bed. For example, General Game Playing (GGP[87] is designed to evaluate AI’s ability in playing different kinds of tabletop games, where rules and states of games are defined with game description language, and has been widely used in AI research and education. Hundreds of single- and multi-player games, such as Go, Tic-Tac-Toe, reversi and minesweeper, are available in GGP. Botzone [88] focuses more on large multi-player games and is famous for its Mahjong AI Competition. Botzone also support 2- and 6-person Texas hold’em. Besides some popular classic games, OpenSpiel [89] provides Poker and Bridge, as well as some classic algorithms and visual tools. Different from GGP and Botzone, Tabletop Games Framework (TAG) [90] is designed to study general AI in modern tabletop games and has implemented more than 20 games, including Uno, Love Letter and Virus!. An extra analysis tool is equipped in TAG for studying action space, branching factors, hidden information, and so on. Ludii [114] is another general game-playing system based on similar principles to Ludi [115] but with different mechanisms. It includes various types of games such as board games, boardless games, stacking games, single- and multi-player games. Ludii uses ludemes as game description grammar. The description language is easy to read and understand according to [114]. Pgx [92] is an environment of board games for RL research. It is written in JAX and can do parallel simulations via GPU/TPU. Pgx can hold single-/multi-agent board games and Atari-like environments including Shogi, Go, Chess, and Backgammon. According to [92], Pgx can perform 10-100 times higher throughput compared to other Python libraries.
@@
OpenAI Gym [93] and PettingZoo [94] offer a number of grid- and pixel-based 2D games. OpenAI Gym [93] provides an API standard for RL research, and contains massive games (e.g., Atari), and a wide range of third-party environments. However, Gym typically focuses on the single-agent setting. To conduct research in multi-agent RL, PettingZoo [94] extends the API standard of Gym and integrates some famous environments like multi-player Atari games, classic card/board games and multi-agent particle environment [116]. As providing the standard API, Gym and PettingZoo are now the most used and compatible platforms, which enhance the evaluation and reproduction in the RL community with a higher speed.
@@
General Video Games AI (GVGAI) [95] platform and GVGAI gym [96] offer a larger number of grid-based single- and two-player games, and have been used in their associated single-player planing, two-player planing, and single-player learning competitions. Users can create new games and levels in GVGAI in an intuitive way using Video Game Description Language (VGDL) [117]. Minigrid [97] is a single-agent platform compatible with OpenAI Gym. The included games are simple in rendering content but can implement complex mechanisms. Melting Pot [98] is a platform with more than 50 diverse games for evaluating the generalisation of multi-agent RL. Stratega [99] evaluates the generalisation of strategy agents with multiple games and levels with an $n$-player turn-based setting. Profiling and logging tools are integrated for analysis. Usually, grid-based platforms are highly scalable in terms of game mechanisms and levels, so researchers can easily design specific tasks to address their research problems with the help of grid-based game platforms. Griddly [100] particularly provides grid-based real-time strategy games with various numbers of players, with a faster running speed and lower memory than Minigrid and GVGAI.
@@
DeepMind Lab [104] acts as a 3D learning environment which is based on some open-source software such as ioquake3. It provides a set of challenging 3D navigation and puzzle-solving tasks for agents to learn. The primary purpose of DeepMind Lab is to act as a working platform for results in AI, especially deep RL. MiniWorld [105] provides a more accessible way to construct 3D environments for RL and robotics research. It is qualified to generate and simulate environments with rooms, doors, mazes and other objects. MiniWorld is simpler than DeepMind Lab and ViZDoom in the view of engineering, so beginners can get started quickly without mastering Unity3D. ML-agents [106] is an open-source platform that allows to create new games in a simpler and more flexible manner with the help of the Unity3D toolkit.
3.4 Guidance for Playing Games
According to the aim, number of agents, observability, language and competition, games and platforms reviewed in this work are categorised in Tabs. 1 and 2. Those games and platforms are often developed for specific research purposes. Readers can choose suitable games and platforms according to the categories. For example, if one focuses on learning-based research like RL [118] and imitation learning [119], games and platforms with ✓under Learning in the column Aim of Tabs. 1 and 2 may be good choices. Examples of related research using those platforms include the work of Mnih el al. [2] using Atari, and Vinyals el al. [15] using StarCraft II. For research on searching algorithms such as Monte-Carlo tree search [120, 121], games and platforms marked in Planning column are good choices. With respect to different scenarios considering the number of agents, column #Agents provides such information. Single-agent scenarios usually meet deterministic, stochastic and configurable environments, denoted with ,   and , respectively. Sparse reward [122, 123] and long-term exploration [124] are common research topics involved with these scenarios. Besides, three multi-agent cases including adversarial, cooperative and mixed scenarios are denoted by “A", “C" and “M", respectively. Research on game theory [111, 112] and multi-agent RL [108, 109, 110] can benefit from the information for choosing suitable testbeds for experiments. Observability of games and platforms can also be checked in Tabs. 1 and 2, which helps with research in partially observable planning [125] and RL [126].
Card/board-related games and platforms, including Blood bowl [52], Sabberstone [56] and Botzone [88], are suitable for research on search, causality and competitive scenarios in multi-agent systems. GVGAI [127], Stratega [99] and Procgen [102] are developed for the purpose of evaluating the generalisation of AI agents. Both GVGAI and Stratega can be used for research on planning and learning, while Procgen specially focuses on RL.
@@ 3.4 Guidance for Playing Games
Besides, the integrated tools differ between games and platforms. For example. in GVGAI, users can easily construct their games and levels via VGDL. Pommerman [66] supports highly self-defined maps. However, game levels in Procgen can only be generated randomly. It is recommended to get started from Gym [93] and PettingZoo [94] for single- and multi-agent RL research since they are the most commonly used platforms. Some platforms such as DeepMind Lab [104] and ML-agent [106] usually are equipped with some visual tools and loggers to analyse the behaviour of agents, but the requirement to learn about the game engines like Unity is crucial.
@@
4 Designing Games by AI
Table 3: 26 environments for designing games by AI. Although Mario-AI, HOG2, PhyDSL, StepMania, OSU! and MiniDungeons are available before 2016, they have been used for testing and evaluating the procedurally generated Super Mario Bro levels in recent years.
@@ 4 Designing Games by AI
Game Tuning (Section 4.1) Parameter FightingICE Java [128]
@@
RaidEnv Python [129]
2D level Science Birds Java [130]
Content Generation (Section 4.2) Mario AI Java [72]
Morai-Maker-Engine C# [131]
@@
Keiki Python [135]
@@
Other (Section 4.3) Code PhyDSL Java [146]
@@
In recent years, game-based platforms have also been used to create new tasks for AI and to study AI creativity, in particular, the ability of creating new contents. Tab. 3 and Fig. 4 categorise the games and platforms used for game design according to research aim including game tuning and content generation, manipulated content type (e.g., parameter, 2D level, 3D level and narrative), programming language and if they have been used by any AI-related competitions. Some resources such as code and dataset are labelled as the other type. Those platforms provide diverse and representative test-beds for discovering the potentials in AI-involved game design.
4.1 Game Tuning